OpenStudy (anonymous):

Write a pseudocode algorithm to compute the product of the first n positive integers. How many multiplications does your algorithm perform?

OpenStudy (anonymous):

Can you explain in your own words, what this is asking you to do? that is how i start all of my psuedocode

OpenStudy (neer2890):

it's like an algorithm for finding the factorial of a number as factorial starts multiplying from 1 to that number itself. so just write an algorithm to find the factorial. Recurrence relation for factorial of a number is f(n)=n*f(n-1)

OpenStudy (anonymous):

Count the number of multiplicative operators: \[1\underbrace{\times}_{\large1}2\underbrace{\times}_{\large2}\cdots\underbrace{\times}_{\large n-2}(n-1)\underbrace{\times}_{\large n-1} n\]